# Functional normalizing flow for statistical inverse problems of partial differential equations

### v1

@@ 5.1 Conditional Functional Normalizing Flow
Figure 14: The comparison of the approximate posterior obtained by CNF-iVI method and the pCN method. The estimated variance and covariance functions obtained by the pCN algorithm are drawn in blue solid line, and by CNF-iVI algorithm the red dashed line. (a): The background truth of $u$. (b): The mean of approximate posterior obtained by CNF-iVI method. (c):The mean of posterior obtained by pCN method. (d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$.
Figures 12, 13, and 14 present qualitative results of the approximate posterior inferred using the trained conditional network $\mathcal{N}_{\lambda}(\bm{v})$ for three different measurement information sets. To further quantitatively assess the effectiveness of the CNF-iVI method, we calculate the re-simulation error:
re-simulation $\begin{split}\textbf{re-simulation error:}&=E_{(\bm{x},\bm{d})}(E_{\nu_{\lambda}(\bm{v}^{(\bm{x},\bm{d})})}\lVert\mathcal{S}\mathcal{G}(u)-\mathcal{S}\mathcal{G}(u_{truth}^{(\bm{x},\bm{d})})\rVert)\\ &\approx\frac{1}{N_{test}N_{samples}}\sum\limits_{i=1}^{N_{test}}\sum\limits_{j=1}^{N_{samples}}\lVert\mathcal{S}\mathcal{G}(u_{ij})-\mathcal{S}\mathcal{G}(u_{truth}^{(\bm{x}_{i},\bm{d}_{i})})\rVert\\ &=0.10834.\end{split}$ (5.1)
where $u_{truth}^{(\bm{x},\bm{d})}$ represents the ground truth and $\bm{v}^{(\bm{x},\bm{d})}$ denotes the vector corresponding to the measurement information $(\bm{x},\bm{d})$. We set $N_{samples}=500$ for the evaluation.
The visual evidence presented in Figures 12, 13, and 14, combined with the quantitative results from the re-simulation error in (5.1), demonstrates the effectiveness of our proposed method in roughly capturing the properties of the true posterior, conditioned on the measurement information $(\bm{x},\bm{d})$. This highlights the efficiency of conditional functional normalizing flow in roughly inferring the posterior distribution.
(a) Truth
(b) CNF-iVI
(c) pCN
@@
Figure 15: The comparison of the approximate posterior obtained by retrained NF-iVI method and the pCN method. The estimated variance and covariance functions obtained by the pCN algorithm are drawn in blue solid line, and by retrained NF-iVI algorithm the red dashed line. (a): The background truth of $u$. (b): The mean of approximate posterior obtained by retrained NF-iVI method. (c):The mean of posterior obtained by pCN method. (d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$
5.2 Further Training for Specific Data
For the inverse problem corresponding to specific measurement information, we can refine the approximate posterior obtained from the trained conditional network $\mathcal{N}_{\lambda}(\bm{v})$ using a retraining method (Algorithm 3). For the test dataset $\{(\bm{x}_{1},\bm{d}_{1}),(\bm{x}_{2},\bm{d}_{2}),\ldots,(\bm{x}_{N_{test}},\bm{d}_{N_{test}})\}$ used in Subsection 5.1, we convert each pair $(\bm{x}_{i},\bm{d}_{i})$ into its corresponding vector $\bm{v}_{i}$. These vectors are then input into the trained neural network $\mathcal{N}_{\lambda}(\bm{v})$ to obtain initial approximations $\{\nu_{\lambda}(\bm{v}_{1}),\nu_{\lambda}(\bm{v}_{2}),\ldots,\nu_{\lambda}(\bm{v}_{N_{test}})\}$ of the posterior distributions. Subsequently, we refine these approximate posteriors by retraining them using Algorithm 3, resulting in a new set of refined posteriors $\{\hat{\nu}_{1},\hat{\nu}_{2},\ldots,\hat{\nu}_{N_{test}}\}$.
We use Adam to retrain the approximate posterior. Following Algorithm 3, the functional Sylvester flow is retrained for 1000 iterations $(K=1000)$ using 30 samples $(N=30)$ in each iteration, with an initial learning rate of $\alpha_{0}=0.001$. The learning rate follows a step decay schedule with a multiplicative factor of $\tau=0.9$ after every 200 iterations.
(a) Truth
(b) CNF-iVI
(c) pCN
diag
(d) Covariance
(e) Covariance
(f) Covariance
Figure 16: The comparison of the approximate posterior obtained by retrained NF-iVI method and the pCN method. The estimated variance and covariance functions obtained by the pCN algorithm are drawn in blue solid line, and by retrained NF-iVI algorithm the red dashed line. (a): The background truth of $u$. (b): The mean of approximate posterior obtained by retrained NF-iVI method. (c):The mean of posterior obtained by pCN method. (d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$
Figures 15f, 16f, and 17f present the results obtained from retraining the initial outputs from the trained conditional functional normalizing flow in Subsection 5.1. These inverse problems correspond to the same scenarios depicted in Figures 12, 13, and 14, respectively. We use the results from the trained conditional functional normalizing flow as initial values and retrain the flow models using Algorithm 3. In subfigure (a) of Figures 15f, 16f, and 17f, we present the true values. Subfigure (b) of Figures 15f, 16f, and 17f depict the mean values obtained by the pCN algorithm, while subfigure (c) of Figures 15f, 16f, and 17f show the mean values of the retrained models. Additionally, we conduct a detailed comparison of the covariance functions obtained by the two methods. We present the covariance function $c(x,y)$ on the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$, $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$, $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$ in subfigures (d), (e), and (f) of Figures 15f, 16f, and 17f, respectively.
(a) Truth
(b) CNF-iVI
(c) pCN
diag
(d) Covariance
(e) Covariance
(f) Covariance
Figure 17: The comparison of the approximate posterior obtained by retrained NF-iVI method and the pCN method. The estimated variance and covariance functions obtained by the pCN algorithm are drawn in blue solid line, and by retrained NF-iVI algorithm the red dashed line. (a): The background truth of $u$. (b): The mean of approximate posterior obtained by retrained NF-iVI method. (c):The mean of posterior obtained by pCN method. (d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$
Furthermore, we provide numerical evidence to support the effectiveness of the retraining process. The re-simulation error of the retrained model is calculated as follows:
re-simulation $\begin{split}\textbf{re-simulation error:}&=E_{(\bm{x},\bm{d})}(E_{\hat{\nu}_{(\bm{x},\bm{d})}}\lVert\mathcal{S}\mathcal{G}(u)-\mathcal{S}\mathcal{G}(u_{truth}^{(\bm{x},\bm{d})})\rVert)\\ &\approx\frac{1}{N_{train}N_{samples}}\sum\limits_{i=1}^{N_{test}}\sum\limits_{j=1}^{N_{samples}}\lVert\mathcal{S}\mathcal{G}(u_{ij})-\mathcal{S}\mathcal{G}(u_{truth}^{(\bm{x}_{i},\bm{d}_{i})})\rVert\\ &=0.02393,\end{split}$ (5.2)
where $u_{truth}^{(\bm{x},\bm{d})}$ represents the ground truth and $\hat{\nu}_{(\bm{x},\bm{d})}$ denotes the approximate posterior obtained from the retrained flow models corresponding to the measurement information $(\bm{x},\bm{d})$. We set $N_{samples}=500$ for the evaluation, similar to (5.1). This demonstrates a significant improvement compared to the re-simulation error obtained without retraining.
@@ 5.2 Further Training for Specific Data
To numerically compare the covariance operator, we also compute the relative errors of the variance functions and the covariance functions between estimated posteriors generated by pCN algorithm and retrained NF-iVI method in Table 13. In the table, the notation $\bm{c}$ means the total relative error, and the notation $\{c(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ means the covariance function values on the pair of mesh points $\{(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ ($k=0,40,80$). The numbers below this notation are the relative errors between the vectors obtained by pCN method and retrained NF-iVI method for three different measurement information, respectively.
The visual evidence presented in Figures 15f, 16f, and 17f, combined with the quantitative results from the re-simulation error in (5.2), demonstrates the effectiveness of Algorithm 3 in refining the posterior approximation obtained from conditional functional normalizing flow.
@@ 6 Conclusion
In this paper, we introduce the NF-iVI and CNF-iVI methods within the infinite-dimensional space, providing an efficient computational framework for applying variational inference to inverse problems in function spaces. NF-iVI constructs a transformation that ensures measure equivalence while offering sufficient flexibility for the transformed measure. Building upon NF-iVI, CNF-iVI incorporates a conditional neural network to control the parameters of functional normalizing flow, significantly reducing computational costs.
The established NF-iVI and CNF-iVI approaches can be applied to PDE inverse problems with Gaussian priors, leading to explicit forms of the approximate posterior measure. We have successfully applied these methods to two inverse problems: the simple smooth equation and the steady-state Darcy flow problem. For all these inverse problems, NF-iVI provides an approximate posterior that closely resembles the ground truth generated by the pCN algorithm. While CNF-iVI may initially provides a rough estimate of the posterior, applying the adjusted Algorithm 3 can significantly improve its accuracy.
The current NF-iVI and CNF-iVI methods are based on transformations such as planar flow, Householder flow, Sylvester flow, and projected transformation flow. Future research could explore other types of neural networks that satisfy the general theorem proposed in infinite-dimensional space. For instance, in Euclidean spaces, the non-centered parameterisations formulation [36] has been employed as a transformation for normalizing flow. In this model, we could introduce different types of neural networks into each step of the transformation and carefully design each layer to enhance its complexity. Therefore, developing the NF-iVI method based on Non-centered parameterisations and combining it with other types of neural networks in function spaces, such as Fourier neural operators [33], is a promising avenue for future research.
@@ 7.1 Functional Determinant
To obtain a precise expression for the Radon-Nikodym derivative between $\mu_{f}$ and $\mu_{0}$ as discussed in Subsection 2.2, we will introduce the concept of the Fredholm-Carleman determinant. This determinant is applicable to operators of the form $I+\mathcal{K}$, where $\mathcal{K}$ is a finite-dimensional operator mapping from a separable Hilbert space $\mathcal{H}_{u}$ to itself, and $I$ denotes the identity operator.
@@
We refer to ${\det}_{1}(I+\mathcal{K})$ as the Fredholm-Carleman determinant associated with the operator $I+\mathcal{K}$, and ${\det}_{2}(I+\mathcal{K})$ as the regularized Fredholm-Carleman determinant of the same operator. For further details, the reader is directed to Chapter 6.4 in the reference [7].
@@
  • Proof.
    The proof can be divided in two steps.
@@
    Step 2: We can easily find that $I+\mathcal{T}$ is Fréchet differentiable and
@@
    Relying on Steps 1 and 2 and Example 10.27 of [8], we obtain that $\mu\sim\mu_{0}$ where $\mu=\mu_{0}\circ(f^{(N)}\circ f^{(N-1)}\circ\cdots\circ f^{(1)})^{-1}$. ∎
7.3 Proof of Theorem 2.4
  • Proof.
    Firstly, we consider the case where $N=1$, and regard $\mathcal{F}^{(1)}$ as $\mathcal{F}.$ Let $\{e_{1},e_{2},\ldots,e_{M}\}$ be a set of standard orthogonal bases of the space $\mathcal{F}(\mathcal{H}_{u})$. From Theorems 5.8.3 and 6.6.7 and Corollary 6.6.8 of [7], we know that
@@
    $\Lambda_{\mathcal{F}}(u):=\left|{\det}_{2}\left(I+D\mathcal{F}(u)\right)\right|\exp\left[\delta\mathcal{F}(u)-\frac{1}{2}\lVert\mathcal{F}(u)\rVert_{\mathcal{H}}^{2}\right],$
@@
    where for any $i\in\{1,2,\ldots,M\}$, $\mathcal{F}_{i}(u)=\langle\mathcal{F}(u),e_{i}\rangle_{\mathcal{H}}$, $\widehat{e}_{i}(u)=\langle u,e_{i}\rangle_{\mathcal{H}}$. By simplifying the calculation, we know that
    $\begin{split}\sum\limits_{i=1}^{M}\widehat{e}_{i}(u)\mathcal{F}_{i}(u)=\left\langle\sum\limits_{i=1}^{M}\widehat{e}_{i}(u)e_{i},\sum\limits_{i=1}^{M}\mathcal{F}_{i}(u)e_{i}\right\rangle_{\mathcal{H}}=\left\langle u,\mathcal{F}(u)\right\rangle_{\mathcal{H}}.\end{split}$
@@
    where $u_{k}=f_{k}\circ\cdots\circ f_{1}(u)$, $k=1,2,\ldots,N-1$. Note that $\mathcal{T}(u)=f(u)-u$, from Corollary 6.6.8 of [7], we have
@@
    $\text{H}:\mathcal{M}\times[0,1]\rightarrow\mathcal{H}_{u}$
@@
    which ensures that $h_{t}(q)=q+\text{H}(q,t)\neq 0$. By the compact homotopy invariance of the Leray-Schauder degree [19], we obtain
@@
      The operator $\mathcal{F}_{\theta_{n}}^{(n)}$ is a finite rank operator.
@@
    Since $\mathcal{A}_{n}$ satisfies $\text{Im}(\mathcal{A}_{n})\subset\mathcal{H}$, we know that $\mathcal{F}_{\theta_{n}}^{(n)}\subset\mathcal{H}$. At the same time, since the domain $D(\mathcal{A}_{n})$ of operator $\mathcal{A}_{n}$ has finite dimensions and $\mathcal{A}_{n}$ is a linear operator, we can conclude that $\text{Im}(\mathcal{A}_{n})$ has finite dimensions, which means that $\mathcal{F}_{\theta_{n}}^{(n)}$ is a finite rank operator. The above discussion verifies the first two conditions required by Theorem 2.4.
    Because $\mathcal{A}_{n}$ is a finite rank linear operator, it must be a bounded operator. Note that $h(x)=tanh(x)\subset(-1,1)$, so $h(\mathcal{B}_{n}u+b_{n})$ is bounded in $\mathbb{R}^{M}$ for any $u\in\mathcal{H}_{u}$. Consequently, $\mathcal{A}_{n}h(\mathcal{B}_{n}u+b_{n})$ is bounded in $\mathcal{H}_{u}$, implies that the range of $\mathcal{F}_{n}$ is bounded in $\mathcal{H}_{u}$. By Lemma 2.2, we know that $I+\mathcal{F}_{\theta_{n}}^{(n)}$ is bijective. This verifies the third condition required by Theorem 2.4.
@@
    $D\mathcal{F}_{\theta_{n}}^{(n)}(u)$ $=\mathcal{A}_{n}Dh(\mathcal{B}_{n}u+b_{n})\mathcal{B}_{n}=\mathcal{A}_{n}\mathrm{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n},$
    and $\mathcal{A}_{n}\text{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}$ has the same point spectrum with $\text{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$ (see Theorem 2.9 for details). Furthermore, we recognize that $\text{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))$ is a diagonal matrix with diagonal elements within the interval $(0,1)$. Therefore, combined with the conditions that the eigenvalues of $\mathcal{B}_{n}\mathcal{A}_{n}$ are not in $(-\infty,1]$, we find that the eigenvalues of $\mathrm{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$ are not in $(-\infty,1]$, so the point spectrum of $\mathcal{A}_{n}\text{diag}(h^{{}^{\prime}}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}$ is not in $(-\infty,1]$. Thus, we have verified the final condition required by Theorem 2.4.
@@
    For the third condition, note that
@@
    and the operator $\mathcal{Q}R_{n}\mathcal{P}$ has the same eigenvalues with the matrix $R_{n}$, the point spectrum of $D\mathcal{F}_{\theta_{n}}^{(n)}(u)$ is not in $(-\infty,1]$. Finally, Lemma 2.1 implies that $I+\mathcal{F}_{\theta_{n}}^{(n)}$ is invertible. Based on the above discussion, we can conclude that $\mathcal{F}_{\theta_{n}}^{(n)}(u)=\mathcal{Q}R_{n}(\mathcal{P}u+b_{n})$ satisfies the conditions of Theorem 2.4, which completes the proof of the theorem. ∎
@@
    $\mathcal{B}:\mathcal{H}_{u}\rightarrow\mathbb{R}^{M},\quad\mathcal{A}:\mathbb{R}^{M}\rightarrow\mathcal{H}_{u}.$

### v2

@@ 5.1 Conditional Functional Normalizing Flow
Figure 14: The comparison of the measures obtained by retrained NF-iVI method and the pCN method. The estimated variance and covariance functions obtained by the pCN algorithm are drawn in blue solid line, and by retrained NF-iVI algorithm the red dashed line. (a): The background truth of $u$. (b): The mean of approximate posterior obtained by retrained NF-iVI method. (c):The mean of posterior obtained by pCN method. (d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$.
5.2 Further Training for Specific Data
For the inverse problem corresponding to a specific measurement pair, we can refine the approximate posterior obtained from the trained conditional network $\mathcal{N}_{\lambda}(\bm{v})$ using a retraining method. Consider the test dataset $\{(\bm{x}_{1},\bm{d}_{1}),(\bm{x}_{2},\bm{d}_{2}),\ldots,(\bm{x}_{N_{test}},\bm{d}_{N_{test}})\}$ used in Subsection 5.1, we first convert each pair $(\bm{x}_{i},\bm{d}_{i})$ into its corresponding vector $\bm{v}_{i}$, and then inputting $\bm{v}_{i}$ into the trained neural network $\mathcal{N}_{\lambda}(\bm{v})$ to obtain initial approximations $\{\nu_{\lambda}(\bm{v}_{1}),\nu_{\lambda}(\bm{v}_{2}),\ldots,\nu_{\lambda}(\bm{v}_{N_{test}})\}$ of the posteriors. Subsequently, we refine these approximate posteriors with Algorithm 1 using $\theta_{0}^{(i)}=\mathcal{N}_{\lambda}(\bm{v}_{i})$ as a start point, resulting in a new set of refined posteriors $\{\hat{\nu}_{1},\hat{\nu}_{2},\ldots,\hat{\nu}_{N_{test}}\}$.
We use Adam to retrain the approximate posterior. Following Algorithm 1, the functional Sylvester flow is retrained for 500 iterations $(K=500)$ using 30 samples $(N=30)$ in each iteration, with an initial learning rate of $\alpha_{0}=0.001$. The learning rate follows a step decay schedule with a multiplicative factor of $\tau=0.9$ after every 100 iterations. Figures 14 and 15 present the refined results for the same scenarios shown in Figures 12 and 13, obtained by retraining the initial output of the conditional flow model from Subsection 5.1.
(a) background truth
(b) mean (retrained model)
(c) mean (pCN)
@@ 5.2 Further Training for Specific Data
Figure 15: The comparison of the measures obtained by retrained NF-iVI method and the pCN method. The estimated variance and covariance functions obtained by the pCN algorithm are drawn in blue solid line, and by retrained NF-iVI algorithm the red dashed line. (a): The background truth of $u$. (b): The mean of approximate posterior obtained by retrained NF-iVI method. (c):The mean of posterior obtained by pCN method. (d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$.
@@
To numerically compare the covariance operators, we also compute the relative errors of the variance functions and covariance functions between the estimated posteriors generated by the pCN algorithm and the retrained NF-iVI method, as reported in Table 5. In the table, the notation $\bm{c}$ denotes the total relative error, while $\{c(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ represents the values of the covariance function evaluated at the pairs of mesh points $\{(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ ($k=0,40,80$). The numerical values reported correspond to the relative errors between the vectors obtained using the pCN method and those obtained using the retrained NF-iVI method, for two different measurement pairs, respectively.
Furthermore, we provide numerical evidence to support the effectiveness of the retraining process. The re-simulation error of the retrained model is computed as
$\begin{split}\textbf{re-simulation error:}&=E_{(\bm{x},\bm{d})}\left(E_{\hat{\nu}_{(\bm{x},\bm{d})}}\lVert\mathcal{S}\mathcal{G}(u)-\mathcal{S}\mathcal{G}(u_{truth}^{(\bm{x},\bm{d})})\rVert\right)\\ &\approx\frac{1}{N_{test}N_{samples}}\sum\limits_{i=1}^{N_{test}}\sum\limits_{j=1}^{N_{samples}}\lVert\mathcal{S}\mathcal{G}(u_{ij})-\mathcal{S}\mathcal{G}(u_{truth}^{(\bm{x}_{i},\bm{d}_{i})})\rVert\\ &=0.02393,\end{split}$ (5.2)
where $u_{truth}^{(\bm{x},\bm{d})}$ represents the ground truth and $\hat{\nu}_{(\bm{x},\bm{d})}$ denotes the approximate posterior obtained from the retrained flow models corresponding to the measurement pair $(\bm{x},\bm{d})$. We set $N_{samples}=500$ for the evaluation, consistent with (5.1). This result demonstrates a significant improvement over the re-simulation error (5.1) obtained without retraining.
The visual evidence presented in Figures 14 and 15, together with the re-simulation error in (5.2) and the relative errors reported in Table 5, demonstrates the effectiveness of the retraining method.
@@ 6 Conclusion
In this paper, we introduce the NF-iVI and CNF-iVI methods in an infinite-dimensional setting, providing an efficient computational framework for applying variational inference to inverse problems in function spaces. The NF-iVI method constructs a transformation that guarantees measure equivalence while retaining sufficient flexibility in the transformed measure. Building on NF-iVI, the CNF-iVI method incorporates a conditional neural network to control the parameters of the functional normalizing flow, thereby significantly reducing computational costs.
The proposed NF-iVI and CNF-iVI approaches are applicable to PDE inverse problems with Gaussian priors and yield explicit representations of the approximate posterior measure. We have successfully applied these methods to three inverse problems: a simple smooth equation, the steady-state Darcy flow problem, and electrical impedance tomography. Although CNF-iVI may initially provide a relatively rough approximation of the posterior distribution, the application of the retraining strategy can substantially improve its accuracy.
The current NF-iVI and CNF-iVI frameworks are based on transformations such as planar flow, Householder flow, Sylvester flow, and projected transformation flow. Future work may explore alternative classes of neural networks that satisfy the general theorem established in infinite-dimensional spaces. For example, in Euclidean settings, the non-centered parameterization formulation papaspiliopoulos2003non has been employed as a transformation within normalizing flows. Under the non-centered parameterization framework, different types of neural networks can be introduced at each stage of the transformation, with carefully designed layers to enhance model expressiveness. Consequently, developing NF-iVI methods with measure-equivalence properties based on non-centered parameterization and integrating them with other neural architectures in function spaces, such as Fourier neural operators li2020fourier, represents a promising direction for future research.
@@ 7.1 Functional Determinant
To obtain a precise expression for the Radon-Nikodym derivative between $\mu_{f}$ and $\mu_{0}$ as discussed in Subsection 2.2, we will introduce the concept of the Fredholm-Carleman determinant. This determinant is applicable to operators of the form $I+\mathcal{K}$, where $\mathcal{K}$ is an operator with finite dimensional range mapping from a separable Hilbert space $\mathcal{H}_{u}$ to itself, and $I$ denotes the identity operator.
@@
We refer to ${\det}_{1}(I+\mathcal{K})$ as the Fredholm-Carleman determinant associated with the operator $I+\mathcal{K}$, and ${\det}_{2}(I+\mathcal{K})$ as the regularized Fredholm-Carleman determinant of the same operator. For further details, the reader is directed to Chapter 6.4 in the reference gaussianmeasure.
@@
We begin by proving the equivalence of measures. The proof can be divided in two steps.
@@
Step 2: We can find that $I+\mathcal{T}$ is Fréchet differentiable and
@@
Relying on Steps 1 and 2 and Example 10.27 of MR2663405, we obtain that $\mu\sim\mu_{0}$ where $\mu=\mu_{0}\circ(f^{(N)}\circ f^{(N-1)}\circ\cdots\circ f^{(1)})^{-1}$.
Next, considering the Radon–Nikodym derivative of $\mu_{f_{\theta}}$ with respect to $\mu_{0}$. We first examine the case $N=1$, identifying $\mathcal{F}^{(1)}$ with $\mathcal{F}$. Let $\{e_{1},e_{2},\ldots,e_{M}\}$ be a set of standard orthogonal bases of the space $\mathcal{F}(\mathcal{H}_{u})\subset\mathcal{H}$. From Corollary 6.6.8 of gaussianmeasure, we know that
@@
$\Lambda_{\mathcal{F}}(u)\mathrel{\mathop{\ordinarycolon}}=\left|{\det}_{2}\left(I+D\mathcal{F}(u)\right)\right|\exp\left[\delta\mathcal{F}(u)-\frac{1}{2}\lVert\mathcal{F}(u)\rVert_{\mathcal{H}}^{2}\right],$
@@
where for any $i\in\{1,2,\ldots,M\}$, $\mathcal{F}_{i}(u)=\langle\mathcal{F}(u),e_{i}\rangle_{\mathcal{H}}$, $\widehat{e}_{i}(u)=\mathcal{C}_{0}^{-1}e_{i}(u)$ ($\widehat{h}$ is called the $\mu_{0}$-measurable linear functional generated by $h$ for any $h\in\mathcal{H}$, detail definition can be found in p.60 of gaussianmeasure). By simplifying the calculation, we denote that
$\begin{split}\left\langle u,\mathcal{F}(u)\right\rangle_{\mathcal{H}}\mathrel{\mathop{\ordinarycolon}}=\sum\limits_{i=1}^{M}\widehat{e}_{i}(u)\mathcal{F}_{i}(u)=\left\langle\sum\limits_{i=1}^{M}\widehat{e}_{i}(u)e_{i},\sum\limits_{i=1}^{M}\mathcal{F}_{i}(u)e_{i}\right\rangle_{\mathcal{H}}.\end{split}$
Remark 7.1.
For $u\sim\mathcal{N}(0,\mathcal{C}_{0})$, the event $u\in\mathcal{H}$ is not guaranteed to occur, and thus the inner product $\left\langle u,\mathcal{F}(u)\right\rangle_{\mathcal{H}}$ is not always well-defined. However, when $u\in\mathcal{H}$, the identity $\left\langle u,\mathcal{F}(u)\right\rangle_{\mathcal{H}}=\sum\limits_{i=1}^{M}\widehat{e}_{i}(u)\mathcal{F}_{i}(u)$ holds. In our numerical experiments, we generate samples from the Gaussian distribution $\mathcal{N}(0,\mathcal{C}_{0})$ by truncating its Karhunen-Loève (KL) expansion. This sampling method ensures that all generated samples $u$ satisfy $u\in\mathcal{H}$. Consequently, the use of $\left\langle u,\mathcal{F}(u)\right\rangle_{\mathcal{H}}=\sum\limits_{i=1}^{M}\widehat{e}_{i}(u)\mathcal{F}_{i}(u)$ is appropriate in the numerical context. To simplify the theorem’s presentation, we uniformly employ the notation $\left\langle u,\mathcal{F}(u)\right\rangle_{\mathcal{H}}$ to represent this term.
@@
where $u_{k}=f_{k}\circ\cdots\circ f_{1}(u)$, $k=1,2,\ldots,N-1$. Note that $\mathcal{T}(u)=f(u)-u$, from Corollary 6.6.8 of gaussianmeasure, we have
@@
$\text{H}\mathrel{\mathop{\ordinarycolon}}\mathcal{M}\times[0,1]\rightarrow\mathcal{H}_{u}$
@@
which ensures that $h_{t}(q)=q+\text{H}(q,t)\neq 0$. By the compact homotopy invariance of the Leray-Schauder degree MR1987179, we obtain
@@
    The operator $\mathcal{F}_{\theta_{n}}^{(n)}$ has finite dimensional range.
@@
Since $\mathcal{A}_{n}$ satisfies $\text{Im}(\mathcal{A}_{n})\subset\mathcal{H}$, we know that $\mathcal{F}_{\theta_{n}}^{(n)}\subset\mathcal{H}$. At the same time, since the domain $D(\mathcal{A}_{n})$ of operator $\mathcal{A}_{n}$ has finite dimensions and $\mathcal{A}_{n}$ is a linear operator, we can conclude that $\text{Im}(\mathcal{A}_{n})$ has finite dimensions, which means that $\mathcal{F}_{\theta_{n}}^{(n)}$ has finite dimensional range. The above discussion verifies the first two conditions required by Theorem 2.3.
Since $\mathcal{A}_{n}$ is a linear operator with finite dimensional range, it must be a bounded operator. Note that $h(x)=\tanh(x)\subset(-1,1)$, so $h(\mathcal{B}_{n}u+b_{n})$ is bounded in $\mathbb{R}^{M}$ for any $u\in\mathcal{H}_{u}$. Consequently, $\mathcal{A}_{n}h(\mathcal{B}_{n}u+b_{n})$ is bounded in $\mathcal{H}_{u}$, implies that the range of $\mathcal{F}_{n}$ is bounded in $\mathcal{H}_{u}$. By Lemma 2.2, we know that $I+\mathcal{F}_{\theta_{n}}^{(n)}$ is bijective. This verifies the third condition required by Theorem 2.3.
@@
$D\mathcal{F}_{\theta_{n}}^{(n)}(u)$ $=\mathcal{A}_{n}Dh(\mathcal{B}_{n}u+b_{n})\mathcal{B}_{n}=\mathcal{A}_{n}\mathrm{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n},$
and $\mathcal{A}_{n}\text{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}$ has the same point spectrum with $\text{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$ (see Theorem 2.8 for details). Furthermore, we recognize that $\text{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))$ is a diagonal matrix with diagonal elements within the interval $(0,1)$. Therefore, combined with the conditions that the eigenvalues of $\mathcal{B}_{n}\mathcal{A}_{n}$ are not in $(-\infty,1]$, we find that the eigenvalues of $\mathrm{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}\mathcal{A}_{n}$ are not in $(-\infty,1]$, so the point spectrum of $\mathcal{A}_{n}\text{diag}(h^{\prime}(\mathcal{B}_{n}u+b_{n}))\mathcal{B}_{n}$ is not in $(-\infty,1]$. Thus, we have verified the final condition required by Theorem 2.3.
@@
For the third and fourth conditions, note that
@@
and the operator $\mathcal{Q}R_{n}\mathcal{P}$ has the same eigenvalues with the matrix $R_{n}$, the point spectrum of $D\mathcal{F}_{\theta_{n}}^{(n)}(u)$ is not in $(-\infty,1]$. Thus, Lemma 2.1 implies that $I+\mathcal{F}_{\theta_{n}}^{(n)}$ is invertible. Based on the above discussion, we can conclude that $\mathcal{F}_{\theta_{n}}^{(n)}(u)=\mathcal{Q}R_{n}(\mathcal{P}u+b_{n})$ satisfies the conditions of Theorem 2.3, which completes the proof of the theorem. ∎
@@
$\mathcal{B}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}_{u}\rightarrow\mathbb{R}^{M},\quad\mathcal{A}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{M}\rightarrow\mathcal{H}_{u}.$